Please help me solve this problem: The difference of nine times a number and 6 results in five times the difference of twice the number and 3.
Found 2 solutions by rothauserc, algebra hello:
Answer by rothauserc(4718)

let x be the number, then
9x -6 = 5(2x -3)
:
9x -6 = 10x -15
:
x = 9
:
to check substitute for x in the first equation
9(9) -6 = 5(2(9) -3)
:
81 -6 = 5(18 -3)
:
75 = 5(15)
:
75 = 75
answer checks

Answer by algebra hello(55)

let the number is x then according to question 9x-6=5(2x-3) 9x-6=10x-15 9x-10x=-15 6 -x=-9 x=9